OpenStudy (anonymous):

Divide these two complex numbers. (2-4i)/(4-3i)

OpenStudy (anonymous):

multiply numerator and denominator by \[4+3i\] because \[(4-3i)(4+3i)=4^2+3^2=25\]

OpenStudy (ksaimouli):

25

OpenStudy (anonymous):

so is it (4-10i)/25

OpenStudy (anonymous):

well 25 is the denominator. the numerator is \[(2-4i)(4+3i)=20-10i\]

OpenStudy (anonymous):

so you get \[\frac{20-10i}{25}=\frac{4-2i}{5}\] or in standard form \[\frac{4}{5}-\frac{2}{5}i\]

OpenStudy (anonymous):

oh alright, thank you so much!